# LLM-Augmented Changepoint Detection: A Framework for Ensemble Detection and Automated Explanation

### v1

@@ A Framework for Ensemble Detection and Automated Explanation
Fabian Lukassen1, Christoph Weisser1, Michael Schlee1, Manish Kumar2, Anton Thielmann1, Benjamin Saefken2, Thomas Kneib1 1University of Göttingen  2TU Clausthal fabian.lukassen@stud.uni-goettingen.de {tkneib, michael.schlee}@uni-goettingen.de christoph.weisser@oxon.org {benjamin.saefken, manish.kumar.2}@tu-clausthal.de antonthielmann@t-online.de
@@ 1 Introduction
Changepoint detection represents a fundamental challenge in time series analysis, with applications spanning economics, environmental science, public health, and numerous other domains truong2020ruptures. In general, changepoint detection involves identifying moments when the statistical properties of a time series undergo an abrupt change, indicating regime changes, policy interventions, market crashes, or other significant events. Although traditional statistical methods have shown effectiveness in detecting the temporal location of these changes, they face two persistent limitations that our work addresses.
First, no single detection algorithm performs optimally across all types of time series and changepoint characteristics. Following the spirit of the no free lunch theorem wolpert1997no, methods such as CUSUM page1954cusum are excellent at detecting mean shifts in stationary series, while others like the Bai-Perron test bai2003computation are superior for multiple breaks in regression settings, and PELT killick2012pelt handles complex change patterns efficiently. This diversity of strengths and weaknesses necessitates careful method selection based on data characteristics, a process that requires substantial expertise and often involves trial-and-error approaches.
Second, traditional methods provide only statistical evidence of the changepoints without contextual explanations of why these changes occurred. Analysts must manually investigate historical records, news archives, and domain-specific knowledge to understand the underlying causes merelo2024pivotal. This interpretive burden is time-consuming, subjective, and increasingly impractical as data volumes grow exponentially across industries and research domains.
@@ 2.1 Ensemble Changepoint Detection
Although the literature on individual changepoint detection algorithms is extensive aminikhanghahi2017survey, ensemble-based methodologies have recently emerged as a means to enhance robustness and reliability. Several principled ensemble frameworks have been proposed, each implementing distinct aggregation strategies and exhibiting specific methodological trade-offs.
katser2021cpde proposed an unsupervised framework (CPDE) that aggregates cost functions from multiple methods before performing a search. zhao2019beast introduced BEAST, a Bayesian model averaging approach that provides probabilistic estimates of changepoints. Other techniques include data-centric ensembles like Ensemble Binary Segmentation (EBS), which applies a single algorithm to multiple data subsamples korkas2022ebs, and feature-centric ensembles like PCA-uCPD, which runs detection on different principal components of multivariate data qin2025pcaucpd. Machine learning-based approaches such as ChangeForest londschien2023changeforest and deep learning methods such as WWAggr stepikin2025wwaggr have also been developed.
Our work contributes to this space by creating a practical, flexible ensemble that aggregates the final outputs of diverse, complete algorithms rather than intermediate scores or features. Methods such as CPDE katser2021cpde that aggregate cost functions and WWAggr stepikin2025wwaggr that leverage complex Wasserstein distance calculations follow elaborate strategies. In contrast, our approach prioritizes interpretability and simplicity.
@@ 2.2 Event Attribution and Automated Explanation
Manual attribution of statistical findings to real-world events is a long-standing practice in analytics. Recently, merelo2024pivotal demonstrated a methodology for applying changepoint analysis to historical time series and manually validating detected points against known historical events. Our work automates this process using LLMs.
The idea of using LLMs for explanation is not new, but its application to time series is nascent. zhang2024laserbeam proposed LASER-BEAM to generate narratives explaining forecasts of stock market volatility, a forward-looking task. In contrast, our work focuses on the post-hoc explanation of detected historical changepoints. Other related works like iPrompt singh2022iprompt and TCube sharma2021tcube focus on generating explanations for general data patterns or narrating time series, but do not specifically target the causal attribution of discrete changepoints.
@@ 2.3 LLMs for Time Series Analysis
Recent research has explored the capabilities of LLMs for various time series tasks. Models like AXIS lan2025axis have been developed for explainable anomaly detection, which is related but distinct from changepoint detection. Anomalies are usually single-point deviations, whereas changepoints are persistent shifts in statistical properties. TsLLM parker2025tsllm and other similar models aim to create a unified foundation model for general time series tasks such as forecasting and classification, but do not specialize in the attribution of historical events of structural breaks.
@@ 2.4 Retrieval-Augmented Generation (RAG)
To address the limitations of internal knowledge of LLMs, especially for recent or private data, we incorporate Retrieval-Augmented Generation (RAG) lewis2020rag; reuter2025gptopic. RAG is well-established, but its use for explaining time series is novel to the knowledge of the authors. RAAD-LLM russellgilbert2025raadllm applied a similar concept to anomaly detection in system logs, but our work is the first to use RAG to provide contextually rich and factually grounded explanations for changepoints in any arbitrary time series by retrieving from a user-provided corpus of documents.
@@ 2.5 Research Gaps and Our Contribution
We identify a clear gap in the literature: no existing framework integrates robust, ensemble-based changepoint detection with automated, LLM-powered historical event attribution. Our primary contribution is to bridge this gap, creating a seamless pipeline from statistical detection to human-readable insight. We automate the manual attribution process demonstrated by merelo2024pivotal and adapt the explanatory power of LLMs from forecasting zhang2024laserbeam and anomaly detection lan2025axis to the specific post-hoc challenge of explaining the changepoint.
@@ 4.1.1 Individual Method Selection
Users with domain expertise can select from ten distinct changepoint detection algorithms, grouped by their underlying approach: Statistical Tests such as CUSUM page1954cusum for the detection of mean shifts; Segmentation Methods including PELT killick2012pelt, Binary Segmentation scott1974binary, Bottom-Up keogh2001bottomup, and Window-based methods for optimal partitioning; Kernel-Based Methods like Kernel CPD harchaoui2007kernel for distribution changes; and Bayesian Methods such as Bayesian Online CPD adams2007bayesian for probabilistic estimates. Each method produces changepoint locations. Note that each detection method applies method-specific preprocessing to the data (e.g., stationarity transformations, missing value handling, outlier treatment) and computes confidence scores for detected changepoints using method-specific formulations; complete details for each are provided in Appendix A.
@@ 4.1.3 Ensemble Method
This approach prioritizes interpretability: analysts can inspect which methods voted for each changepoint, understand the consensus level, and adjust $v_{\text{min}}$ based on their risk tolerance. Unlike complex aggregation schemes (e.g., Wasserstein barycenters stepikin2025wwaggr), our method is transparent. The complete ensemble procedure is formalized in Algorithm 2.
@@ 4.3 RAG for Private Data
We use Sentence-Transformers reimers2019sentencebert with the all-MiniLM-L6-v2 model (384 dimensions, 80MB) as the default. For higher quality, users can opt for all-mpnet-base-v2 (768 dimensions, 420MB). Given a document chunk $d$, the embedding function is:
@@
We use ChromaDB chromadb2023 as the persistent vector database. Documents are stored with metadata (date, title, type) for hybrid search.
@@ 5.1 Ensemble Method
The datasets represent different types of real-world changepoints: Nile (annual river flow, 1871–1970) shows a mean shift when the Aswan Low Dam construction began in 1898 cobb1978nile; Seatbelts (monthly UK road casualties, 1976–1984) exhibits a sudden drop following the 1983 compulsory seatbelt law harvey1986seatbelt; LGA (monthly LaGuardia Airport passengers, 1977–2015) captures the immediate and sustained impact of the September 11 attacks on air travel ito2005airline911; Ireland Debt (annual debt-to-GDP ratio, 2000–2020) shows the dramatic surge following the 2008 banking crisis and subsequent bailout lane2011irishcrisis; Ozone (annual Antarctic ozone measurements, 1961–2014) marks the reversal point when Montreal Protocol effects began solomon2016ozone; Robocalls (monthly US call volume, 2015–2019) increases sharply after a 2018 federal court ruling loosened FCC restrictions; and Japan Nuclear (annual nuclear share of electricity, 1985–2024) drops precipitously after the 2011 Fukushima disaster led to reactor shutdowns hayashi2013fukushima. Most datasets are sourced from the Turing Change Point Dataset van2020evaluation.
@@
For event attribution, we query three LLMs spanning different scales and architectures: Llama-3.1-8B grattafiori2024llama, a small open-source model run locally; GPT-4o openai2024gpt4o, a large commercial model; and DeepSeek-R1 guo2025deepseekr1, a 671B-parameter Mixture-of-Experts model with 37B active parameters, optimized for reasoning via reinforcement learning. This selection examines how model scale and reasoning specialization affect explanation quality.
@@
Detection is considered correct if the predicted changepoint falls within $\pm$3 data points of the ground truth index. For evaluating LLM-generated explanations, we employ the LLM-as-a-judge framework zheng2023llmjudge. Specifically, we use Kimi K2 kimik2 as the judge. The judge receives the LLM’s explanation along with the ground truth event cause and is asked to determine whether the explanation correctly identifies the underlying event responsible for the changepoint. The judge outputs a binary correctness label (correct/incorrect) based on whether the core causal event is accurately attributed. The exact evaluation prompt is provided in Appendix C.
@@ 7 Conclusion
Confidence scores at both levels are heuristic rather than statistically rigorous. At the individual method level, confidence calculations vary: some derive from test statistics scaled by critical values (CUSUM), others from p-value transformations (Bai-Perron: $1-p_{\text{value}}$), and others from heuristics like local variance reduction (Binary Segmentation). These heterogeneous measures are not calibrated probabilities. At the ensemble level, the voting-based aggregation reflects method agreement rather than true probabilistic uncertainty. Unlike Bayesian approaches such as BEAST zhao2019beast that yield principled posterior distributions, our metrics prioritize interpretability and computational efficiency over statistical formalism. See Appendix A for detailed confidence calculations per method.
@@ A.2 Changepoint Methods (Ruptures-based)
These methods use the ruptures package truong2020ruptures with custom preprocessing and confidence calculations.

### v3

@@ A Framework for Ensemble Detection and Automated Explanation
Fabian Lukassen1, Christoph Weisser3, Michael Schlee1, Manish Kumar2, Anton Thielmann1thanks: Work completed while at TU Clausthal., Benjamin Saefken2, Alexander Silbersdorff1, Thomas Kneib1 1University of Göttingen  2TU Clausthal  3 Hochschule Bielefeld fabian.lukassen@stud.uni-goettingen.de {tkneib, asilbersdorff, michael.schlee}@uni-goettingen.de christoph.weisser@hsbi.de {benjamin.saefken, manish.kumar.2}@tu-clausthal.de antonthielmann@t-online.de
@@ 1 Introduction
Changepoint detection represents a fundamental challenge in time series analysis, with applications spanning economics, environmental science, public health, and numerous other domains Truong et al. (2020). In general, changepoint detection involves identifying moments when the statistical properties of a time series undergo an abrupt change, indicating regime changes, policy interventions, market crashes, or other significant events. Although traditional statistical methods have shown effectiveness in detecting the temporal location of these changes, they face two persistent limitations that our work addresses.
First, no single detection algorithm performs optimally across all types of time series and changepoint characteristics. Following the spirit of the no free lunch theorem Wolpert and Macready (1997), methods such as CUSUM Page (1954) are excellent at detecting mean shifts in stationary series, while others like the Bai-Perron test Bai and Perron (2003) are superior for multiple breaks in regression settings, and PELT Killick et al. (2012) handles complex change patterns efficiently. This diversity of strengths and weaknesses necessitates careful method selection based on data characteristics, a process that requires substantial expertise and often involves trial-and-error approaches.
Second, traditional methods provide only statistical evidence of the changepoints without contextual explanations of why these changes occurred. Analysts must manually investigate historical records, news archives, and domain-specific knowledge to understand the underlying causes Merelo-Guervós (2024). This interpretive burden is time-consuming, subjective, and increasingly impractical as data volumes grow exponentially across industries and research domains.
@@ 2.1 Ensemble Changepoint Detection
Although the literature on individual changepoint detection algorithms is extensive Aminikhanghahi and Cook (2017), ensemble-based methodologies have recently emerged as a means to enhance robustness and reliability. Several principled ensemble frameworks have been proposed, each implementing distinct aggregation strategies and exhibiting specific methodological trade-offs.
Katser et al. (2021) proposed an unsupervised framework (CPDE) that aggregates cost functions from multiple methods before performing a search. Zhao et al. (2019) introduced BEAST, a Bayesian model averaging approach that provides probabilistic estimates of changepoints. Other techniques include data-centric ensembles like Ensemble Binary Segmentation (EBS), which applies a single algorithm to multiple data subsamples Korkas (2022), and feature-centric ensembles like PCA-uCPD, which runs detection on different principal components of multivariate data Qin et al. (2025). Machine learning-based approaches such as ChangeForest Londschien et al. (2023) and deep learning methods such as WWAggr Stepikin et al. (2025) have also been developed.
Our work contributes to this space by creating a practical, flexible ensemble that aggregates the final outputs of diverse, complete algorithms rather than intermediate scores or features. Methods such as CPDE Katser et al. (2021) that aggregate cost functions and WWAggr Stepikin et al. (2025) that leverage complex Wasserstein distance calculations follow elaborate strategies. In contrast, our approach prioritizes interpretability and simplicity.
@@ 2.2 Event Attribution and Automated Explanation
Manual attribution of statistical findings to real-world events is a long-standing practice in analytics. Recently, Merelo-Guervós (2024) demonstrated a methodology for applying changepoint analysis to historical time series and manually validating detected points against known historical events. Our work automates this process using LLMs.
The idea of using LLMs for explanation is not new, but its application to time series is nascent. Zhang et al. (2024) proposed LASER-BEAM to generate narratives explaining forecasts of stock market volatility, a forward-looking task. In contrast, our work focuses on the post-hoc explanation of detected historical changepoints. Other related works like iPrompt Singh et al. (2023) and TCube Sharma et al. (2021) focus on generating explanations for general data patterns or narrating time series, but do not specifically target the causal attribution of discrete changepoints.
@@ 2.3 LLMs for Time Series Analysis
Recent research has explored the capabilities of LLMs for various time series tasks. Models like AXIS Lan et al. (2025) have been developed for explainable anomaly detection, which is related but distinct from changepoint detection. Anomalies are usually single-point deviations, whereas changepoints are persistent shifts in statistical properties. TsLLM Parker et al. (2025) and other similar models aim to create a unified foundation model for general time series tasks such as forecasting and classification, but do not specialize in the attribution of historical events of structural breaks.
@@ 2.4 Retrieval-Augmented Generation (RAG)
To address the limitations of internal knowledge of LLMs, especially for recent or private data, we incorporate Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Reuter et al. (2025). RAG is well-established, but its use for explaining time series is novel to the knowledge of the authors. RAAD-LLM Russell-Gilbert et al. (2025) applied a similar concept to anomaly detection in system logs, but our work is the first to use RAG to provide contextually rich and factually grounded explanations for changepoints in any arbitrary time series by retrieving from a user-provided corpus of documents.
@@ 2.5 Research Gaps and Our Contribution
We identify a clear gap in the literature: no existing framework integrates robust, ensemble-based changepoint detection with automated, LLM-powered historical event attribution. Our primary contribution is to bridge this gap, creating a seamless pipeline from statistical detection to human-readable insight. We automate the manual attribution process demonstrated by Merelo-Guervós (2024) and adapt the explanatory power of LLMs from forecasting Zhang et al. (2024) and anomaly detection Lan et al. (2025) to the specific post-hoc challenge of explaining the changepoint.
@@ 4.1.1 Individual Method Selection
Users with domain expertise can select from ten distinct changepoint detection algorithms, grouped by their underlying approach: Statistical Tests such as CUSUM Page (1954) for the detection of mean shifts; Segmentation Methods including PELT Killick et al. (2012), Binary Segmentation Scott and Knott (1974), Bottom-Up Keogh et al. (2001), and Window-based methods for optimal partitioning; Kernel-Based Methods like Kernel CPD Harchaoui and Cappé (2007) for distribution changes; and Bayesian Methods such as Bayesian Online CPD Adams and MacKay (2007) for probabilistic estimates. Each method produces changepoint locations. Note that each detection method applies method-specific preprocessing to the data (e.g., stationarity transformations, missing value handling, outlier treatment) and computes confidence scores for detected changepoints using method-specific formulations; complete details for each are provided in Appendix A.
@@ 4.1.3 Ensemble Method
This approach prioritizes interpretability: analysts can inspect which methods voted for each changepoint, understand the consensus level, and adjust $v_{\text{min}}$ based on their risk tolerance. Unlike complex aggregation schemes (e.g., Wasserstein barycenters Stepikin et al. (2025)), our method is transparent. The complete ensemble procedure is formalized in Algorithm 2.
@@ 4.3 RAG for Private Data
We use Sentence-Transformers Reimers and Gurevych (2019) with the all-MiniLM-L6-v2 model (384 dimensions, 80MB) as the default. For higher quality, users can opt for all-mpnet-base-v2 (768 dimensions, 420MB). Given a document chunk $d$, the embedding function is:
@@
We use ChromaDB Chroma (2023) as the persistent vector database. Documents are stored with metadata (date, title, type) for hybrid search.
@@ 5.1 Ensemble Method
The datasets represent different types of real-world changepoints: Nile (annual river flow, 1871–1970) shows a mean shift when the Aswan Low Dam construction began in 1898 Cobb (1978); Seatbelts (monthly UK road casualties, 1976–1984) exhibits a sudden drop following the 1983 compulsory seatbelt law Harvey and Durbin (1986); LGA (monthly LaGuardia Airport passengers, 1977–2015) captures the immediate and sustained impact of the September 11 attacks on air travel Ito and Lee (2005); Ireland Debt (annual debt-to-GDP ratio, 2000–2020) shows the dramatic surge following the 2008 banking crisis and subsequent bailout Lane (2011); Ozone (annual Antarctic ozone measurements, 1961–2014) marks the reversal point when Montreal Protocol effects began Solomon et al. (2016); Robocalls (monthly US call volume, 2015–2019) increases sharply after a 2018 federal court ruling loosened FCC restrictions; and Japan Nuclear (annual nuclear share of electricity, 1985–2024) drops precipitously after the 2011 Fukushima disaster led to reactor shutdowns Hayashi and Hughes (2013). Most datasets are sourced from the Turing Change Point Dataset van den Burg and Williams (2020).
@@
For event attribution, we query three LLMs spanning different scales and architectures: Llama-3.1-8B Grattafiori and others (2024), a small open-source model run locally; GPT-4o OpenAI (2024), a large commercial model; and DeepSeek-R1 Guo et al. (2025), a 671B-parameter Mixture-of-Experts model with 37B active parameters, optimized for reasoning via reinforcement learning. This selection examines how model scale and reasoning specialization affect explanation quality.
@@
Detection is considered correct if the predicted changepoint falls within $\pm$3 data points of the ground truth index. For evaluating LLM-generated explanations, we employ the LLM-as-a-judge framework Zheng et al. (2023). Specifically, we use Kimi K2 Moonshot AI (2025) as the judge. The judge receives the LLM’s explanation along with the ground truth event cause and is asked to determine whether the explanation correctly identifies the underlying event responsible for the changepoint. The judge outputs a binary correctness label (correct/incorrect) based on whether the core causal event is accurately attributed. The exact evaluation prompt is provided in Appendix C.
@@ 7 Conclusion
Confidence scores at both levels are heuristic rather than statistically rigorous. At the individual method level, confidence calculations vary: some derive from test statistics scaled by critical values (CUSUM), others from p-value transformations (Bai-Perron: $1-p_{\text{value}}$), and others from heuristics like local variance reduction (Binary Segmentation). These heterogeneous measures are not calibrated probabilities. At the ensemble level, the voting-based aggregation reflects method agreement rather than true probabilistic uncertainty. Unlike Bayesian approaches such as BEAST Zhao et al. (2019) that yield principled posterior distributions, our metrics prioritize interpretability and computational efficiency over statistical formalism. See Appendix A for detailed confidence calculations per method.
@@ A.2 Changepoint Methods (Ruptures-based)
These methods use the ruptures package Truong et al. (2020) with custom preprocessing and confidence calculations.